OpenStudy (anonymous):

√(75)/2 + ((2√3)/4) whats the answer without decimals like with the radicals?

OpenStudy (anonymous):

This should help: sqrt(75) = sqrt(25*3) = 5 sqrt(3)

OpenStudy (ranga):

does the square root extend to just 75 or 75/2 and just 3 or 3/4?

OpenStudy (anonymous):

its just square root 75 and its over 2. So then would the anwser be 12√3/4?

OpenStudy (ranga):

sqrt(75) / 2 + 2sqrt(3) / 4 = sqrt(25*3) / 2 + sqrt(3) / 2 = 5*sqrt(3) / 2 + sqrt(3) / 2 = sqrt(3) / 2 * (5 + 1) = sqrt(3) / 2 * 6 = 3sqrt(3)

OpenStudy (anonymous):

Thank you!

OpenStudy (ranga):

You are welcome.